A house is 43.0 ft long and 28 ft wide, and has 8.0-ft-high ceilings. What is the volume of the interior of the house in cubic meters and cubic centimeters? cm Need Help? Read ItMaster It
Answer #1

43 ft * .305 m/ft = 13.115 m
28 ft = 8.54 m

8 ft = 2.44m


V = 2.44*8.54*13.115 = 273.28 m^3

Then 1 m = 100 cm

So,

1 m^3 = 100^3 cm^3 = 10^6 cm^3

So,

V = 273.28*10^6 cm^3 Answer

I hope help you !!